### v2

# Reinforcement Learning with Verifiable Rewards: GRPO’s Effective Loss, Dynamics, and Success Amplification

Abstract.
Group Relative Policy Optimization (GRPO) was introduced in (Shao et al., 2024) and used successfully to train DeepSeek-R1 (Guo et al., 2025) models for promoting reasoning capabilities of LLMs using verifiable or binary rewards. We show in this paper that GRPO with verifiable rewards can be written as a Kullback–Leibler ($\mathsf{KL}$) regularized contrastive loss, where the contrastive samples are synthetic data sampled from the old policy. The optimal GRPO policy $\pi_{n}$ can be expressed explicitly in terms of the binary reward, as well as the first- and second-order statistics of the old policy ($\pi_{n-1}$) and the reference policy $\pi_{0}$. Iterating this scheme, we obtain a sequence of policies $\pi_{n}$ for which we can quantify the probability of success $p_{n}$. We show that the probability of success of the policy satisfies a recurrence that converges to a fixed point of a function that depends on the initial probability of success $p_{0}$ and the regularization parameter $\beta$ of the $\mathsf{KL}$ regularizer. We show that the fixed point $p^{*}$ is guaranteed to be larger than $p_{0}$, thereby demonstrating that GRPO effectively amplifies the probability of success of the policy.
1. Introduction
In Reinforcement Learning (RL), a policy is learned by maximizing a reward that encodes constraints or an objective we want the policy to conform to or achieve. Policy gradient methods and actor-critic methods (Sutton and Barto, 1998), enable RL-based training of parametric policies, including Large Language Models (LLMs), particularly when dealing with non-differentiable rewards. Unlike supervised learning or preference optimization, which require labeled training data, reinforcement learning generates synthetic data sampled online from the learned policy as training progresses.
Proximal Policy Optimization (PPO), introduced in (Schulman et al., 2017), is a widely used algorithm that facilitates such training. PPO relies on importance sampling from the model’s previous (“old”) policy while ensuring that updates remain within a certain proximity to the old policy. Policy gradient methods are known for their high variance, and PPO mitigates this by learning a critic that reduces the variance of gradient estimates. The critic normalizes the reward, and PPO’s advantage function—defined as the difference between the reward and the critic’s evaluation—drives the optimization process.
Group Relative Policy Optimization (GRPO) was recently introduced in DeepSeekMath (Shao et al., 2024). GRPO closely follows PPO’s optimization framework but differs in how the advantage is computed. Specifically, GRPO estimates the advantage using Monte Carlo rollouts rather than a learned critic. Additionally, GRPO applies whitening to the advantage function, meaning it standardizes the reward’s mean and variance. These statistics are estimated from a “group” of Monte Carlo rollouts corresponding to samples from the LLM policy conditioned on a single input or query to the policy. Whitening the advantage function has been recognized in many PPO implementations as an important ingredient for training stability (Engstrom et al., 2020; Huang et al., 2024). The main novelty in GRPO lies in computing this whitening using Monte Carlo rollouts conditioned on a single input or prompt in the context of LLM training.
GRPO therefore eliminates the need for training a separate critic network alongside the LLM policy, instead leveraging efficient sampling from the LLM’s policy. This is made feasible by optimized model serving through VLLM (Kwon et al., 2023). GRPO has been employed in the DeepSeek model series, including DeepSeek-v3 (Liu et al., 2024) and DeepSeek-R1 (Guo et al., 2025). DeepSeek-R1 unlocked reasoning capabilities in open-source models, and its success can be attributed to several factors and innovations, among them: (1) A strong pre-trained model (DeepSeek-v3), (2) The reasoning chain of thoughts <think>...<think> <answer>...<answer> and (3) The use of verifiable binary rewards with GRPO to fine-tune the models on reasoning and math tasks.
We focus in this paper on Reinforcement Learning with Verifiable Rewards (RLVR) using GRPO, as recently termed by (Lambert et al., 2024). Following (Lambert et al., 2024), we distinguish three types of verifiable rewards in the context of LLM training:
  1. (1)
    Correctness Verification. This corresponds to a binary reward that can be obtained via string matching between the generated response and a gold-standard answer—if such an answer exists—for example, in math problems with known solutions. This type of reward has been used in (Guo et al., 2025) and subsequently in open-source implementations such as Open-R1 (Hugging Face, 2024) and DeepScaleR (Luo et al., 2025). When a gold-standard answer does not exist, one can resort to an LLM as a judge to assess the correctness of the response within the training loop, as done in deliberative alignment (Guan et al., 2025).
  2. (2)
    Verification via Execution. In code generation, a code interpreter is used to execute the generated code, producing a 0/1 reward for fail/pass. A battery of unit tests can also be executed to verify the correctness of the code, resulting in a binary reward. Open-R1 (Hugging Face, 2024) recently open-sourced this type of reward evaluation.
  3. (3)
    Verifiable Constraints. Finally, formatting constraints on the output or refusals to answer can be enforced using simple binary rewards to guide RL training for LLMs (Guo et al., 2025) (Lambert et al., 2024).
Verifiable reward balance simplicity and bias and are thought to be less prone to reward hacking than reward models learned from preference data . Reward hacking is a common issue in reinforcement learning where the policy learns to over-optimize a reward leading to a lower quality of the model (Gao et al., 2023). While verifiable rewards are more resilient to reward hacking, Lambert et al. (2024) showed that for low regularization of the $\mathsf{KL}$ constraint to the reference model, reward hacking occurs when using verifiable constraints. Hence we study in this paper $\mathsf{KL}$-regularized Reinforcement Learning with Verifiable Rewards using GRPO.
Our main contributions are :
  1. (1)
    We show in Section 2 that GRPO with verifiable Rewards can be cast as an adaptive weighted Contrastive Loss between samples from the old policy with 0/1 rewards.
  2. (2)
    Armed with this contrastive loss interpretation, we show in Section 3 that GRPO dynamics as the old policy is replaced with the current optimal policy, result in a closed form recursion for the optimal policy, where the optimal policy $\pi_{n}$ can be expressed in terms of the reference policy $\pi_{\mathrm{ref}}$, the old policy $\pi_{n-1}$ and the probability of success of the old policy $p_{n-1}$ (the frequency of reward “1” of generated responses for a given prompt).
  3. (3)
    Computing the probability of success under $\pi_{n}$, we show in Section 3 that it satisfies a recursion in time, leading to a fixed point equation. We show in Section 4 that under mild assumptions GRPO’s probability success converges asymptotically to this fixed point solution.
  4. (4)
    We show in Section 4 that this fixed point probability of success is guaranteed to be larger than the probability of success of the reference model, proving thereby that GRPO indeed amplifies the probability of success as observed experimentally.
  5. (5)
    Finally we show in Section 5 that for approximate policies obtained for instance by gradient descent, the probability of success remains close to the fixed point probability of success as long as the approximation, statistical and optimization errors remain small.
2. GRPO With verifiable Rewards as an Adaptive Weighted Contrastive Loss
Let $\rho_{\mathcal{Q}}$ be a distribution of prompts or questions, and let $r$ be a reward function that evaluates the output $o\in\mathcal{O}$ of a policy. As discussed in the introduction, we restrict our analysis to verifiable rewards, meaning binary rewards, $r:\mathcal{Q}\times\mathcal{O}\to\{0,1\}$. Given a prompt $q\sim\rho_{Q}$, let $\pi_{\theta}(o|q)$ be the policy of an LLM, where $o$ represents the sequence outcome and $\theta\in\Theta$ the parameters of the model. $\pi_{\theta_{\text{old}}}$ denotes the “old” policy or the policy from a previous iteration. $\pi_{\mathrm{ref}}$ corresponds to the reference policy, and $\mathsf{KL}$ is the Kullback–Leibler divergence :
$\mathsf{KL}(\pi||\pi_{\mathrm{ref}})=\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\mathbb{E}_{o\sim\pi(.|q)}\log\left(\frac{\pi(o|q)}{\pi_{\text{ref}}(o|q)}\right)$
For a regularization parameter $\beta>0$, we start by recalling GRPO’s optimization problem (Shao et al., 2024) :
$\max_{\theta}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)-\beta\mathsf{KL}(\pi_{\theta}||\pi_{\mathrm{ref}})$ (GRPO-Clip)
where the advantage for an outcome $o$, $A(q,o)$ is given by:
$A(q,o)=\frac{r(q,o)-\mathbb{E}_{o^{\prime}\sim\pi_{\theta_{\text{old}}}(.|q)}r(q,o^{\prime})}{\sqrt{\mathsf{Var}_{o^{\prime}\sim\pi_{\theta_{\text{old}}}(.|q)}r(q,o^{\prime}))}},$ (1)
where $\mathrm{Var}$ is the variance and for $\epsilon\in[0,1]$
$f_{\epsilon}(x,y)=\min(xy,\text{clip}(x,1-\epsilon,1+\epsilon)y).$
To simplify Equation GRPO-Clip, let us consider this objective without the clipping ($\epsilon\to+\infty$) we obtain:
$\max_{\theta}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)}A(q,o)-\beta\mathsf{KL}(\pi_{\theta}||\pi_{\mathrm{ref}})$ (GRPO)
We see that GRPO optimizes the whitened reward (the advantage $A(q,o)$) using importance sampling from the “old” policy while maintaining the optimized policy close to $\pi_{\mathrm{ref}}$ as measured by the $\mathsf{KL}$ divergence. If furthermore the clipping is used as in (GRPO-Clip), the likelihood ratio between the policy and the old policy is maintained within a range $[1-\epsilon,1+\epsilon]$.
2.1. GRPO with Clipping
Note that in our context $x=\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)}>0$ and the advantage $A(q,o)$ can be positive or negative and hence if $A(q,o)>0$ we have :
$f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)$ $=A(q,o)\min\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},\text{clip}(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon,1+\epsilon)\right)$
$=A(q,o)\min\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1+\epsilon\right)$
and if $A(q,o)<0$
$f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)$ $=A(q,o)\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},\text{clip}(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon,1+\epsilon)\right)$
$=A(q,o)\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon\right)$
Recall that our reward $r$ is a verifiable reward that evaluates correctness of a reasoning or the execution of the code meaning that $r(q,o)\in\{0,1\}$. We note the probability of success of the old policy $\pi_{\text{old}}$:
$p:=p_{\theta_{\text{old}}}(q)=\mathbb{P}_{o\sim\pi_{\theta_{\text{old}}(.|q)}}(r(q,o)=1)$ (2)
Hence we have for mean and variance of a Bernoulli random variable :
$\mathbb{E}_{o^{\prime}\sim\pi_{\theta_{\text{old}}}(.|q)}r(q,o^{\prime})=p\text{ and }\mathsf{Var}_{o^{\prime}\sim\pi_{\theta_{\text{old}}}(.|q)}r(q,o^{\prime})=p(1-p).$
Assuming $10<p<$ 1 and replacing mean and variance in the advantage function (1) we obtain :
$A(q,o)=\begin{cases}\frac{1-p}{\sqrt{p(1-p)}}&\text{if }r(q,o)=1,\\ -\frac{p}{\sqrt{p(1-p)}}&\text{if }r(q,o)=0.\end{cases}$
which simplifies to :
$A(q,o)=\begin{cases}\sqrt{\frac{1-p}{p}}&\text{if }r(q,o)=1,\\ -\sqrt{\frac{p}{(1-p)}}&\text{if }r(q,o)=0.\end{cases}$
Figure 1. Weighting of GRPO with the probability of success of the old policy.
Hence we have conditioning on $q$:
$\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)$
$=\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)\mathbbm{1}_{r(q,o)=1}+\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)\mathbbm{1}_{r(q,o)=0}$
$=\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},\sqrt{\frac{1-p}{p}}\right)\mathbbm{1}_{r(q,o)=1}+\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},-\sqrt{\frac{p}{(1-p)}}\right)\mathbbm{1}_{r(q,o)=0}$
$=\sqrt{\frac{1-p}{p}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q),r(q,o)=1}\min\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1+\epsilon\right)-\sqrt{\frac{p}{(1-p)}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q),r(q,o)=0}\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon\right)$
and hence the overall cost is obtained by taking expectation over $q$, note that $p=p_{\theta_{\text{old}}}(q)$:
${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\sqrt{\frac{1-p_{\theta_{\text{old}}}(q)}{p_{\theta_{\text{old}}}(q)}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\min\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1+\epsilon\right)\mathbbm{1}_{r(q,o)=1}}$
$-{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\sqrt{\frac{p_{\theta_{\text{old}}}(q)}{(1-p_{\theta_{\text{old}}}(q))}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon\right)\mathbbm{1}_{r(q,o)=0}}-\beta\mathsf{KL}(\pi_{\theta}||\pi_{\mathrm{ref}})$
We see that GRPO is effectively a weighted contrastive loss that is weighted by ratio depending on the probability of succes of $\pi_{\theta_{\text{old}}}(.|q)$. We see from the weights plots that :
  • if the success probability of old policy is high ($p>0.5$), the weighting for points with success is low since the old policy is already good, and for failing point the weight is high and hence they are more penalized.
  • if the success probability of old policy is low ($p<0.5$), the weighting for points with success is high since we want to reinforce those successes, and for failing points these are still penalized but with a small weight.
Moreover due to the clipping we have:
  • for correct outputs the cost is constant $(1+\epsilon)$ if $\pi_{\theta}(o|q)\geq(1+\epsilon)\pi_{\theta_{\text{old}}}(o|q)$
  • for wrong outputs the cost is $(1-\epsilon)$ if $\pi_{\theta}(o|q)\leq(1-\epsilon)\pi_{\theta_{\text{old}}}(o|q)$,
In summary, the standardized reward or the advantage function used in GRPO results in an interesting adaptive weighted contrastive loss : if the probability of success of the old policy is high, the wrong answers are more penalized than the correct ones are reinforced. If the probability of success of old policy is low, the correct answers are more reinforced than the wrong answers are penalized.
2.2. Stabilized GRPO with Clipping
Note that in the previous section we assumed that $10<p<$ 1, we alleviate this in the following by adding a smoothing factor $\varepsilon\in(0,1]$ in the advantage as follows:
$A(q,o)=\frac{r(q,o)-\mu}{\sqrt{\sigma^{2}+\varepsilon}}.$
This results with the following advantage:
$A(q,o)=\begin{cases}\frac{1-p}{\sqrt{p(1-p)}+\epsilon}&\text{if }r(q,o)=1,\\ -\frac{p}{\sqrt{p(1-p)}+\epsilon}&\text{if }r(q,o)=0.\end{cases}$ (3)
Let us denote
$\omega^{+}_{\varepsilon}(p)$ $=\frac{1-p}{\sqrt{p(1-p)+\varepsilon}}$
$\omega^{-}_{\varepsilon}(p)$ $=\frac{p}{\sqrt{p(1-p)+\varepsilon}}.$ (4)
Replacing the stabilized advantage in Equation (GRPO-Clip), conditionally on a prompt $q$ we obtain the following contrastive loss:
$\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)\mathbbm{1}_{r(q,o)=1}+\ \mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)\mathbbm{1}_{r(q,o)=0}$
$=\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},\frac{1-p}{\sqrt{p(1-p)+\varepsilon}}\right)\mathbbm{1}_{r(q,o)=1}+\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},\frac{-p}{\sqrt{p(1-p)+\varepsilon}}\right)\mathbbm{1}_{r(q,o)=0}$
$=\omega^{+}_{\varepsilon}(p)\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\min\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1+\epsilon\right)\mathbbm{1}_{r(q,o)=1}-\omega^{-}_{\varepsilon}(p)\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon\right)\mathbbm{1}_{r(q,o)=0},$
which results in the following contrastive optimization :
$\max_{\theta}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\Big{\{}\omega^{+}_{\varepsilon}(p_{\theta_{\text{old}}}(q))\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\min\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1+\epsilon\right)\mathbbm{1}_{r(q,o)=1}...$
.$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}....-\omega^{-}_{\varepsilon}(p_{\theta_{\text{old}}}(q))\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon\right)\mathbbm{1}_{r(q,o)=0}\Big{\}}-\beta\mathrm{KL}(\pi_{\theta}||\pi_{\mathrm{ref}})$
2.3. Stabilized GRPO with No Clipping
Taking the clipping parameter $\epsilon\to\infty$ we obtain GRPO with no clipping equivalent contrastive optimization as follows:
$\max_{\theta}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\Big{\{}\omega^{+}_{\varepsilon}(p_{\theta_{\text{old}}}(q))\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)}\mathbbm{1}_{r(q,o)=1}..$
..$~{}~{}..-\omega^{-}_{\varepsilon}(p_{\theta_{\text{old}}}(q))\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)}\mathbbm{1}_{r(q,o)=0}\Big{\}}-\beta\mathrm{KL}(\pi_{\theta}||\pi_{\mathrm{ref}})$ (GRPO-No-Clip)
which is equivalent to the following problem:
$\max_{\theta}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\Big{\{}\omega^{+}_{\varepsilon}(p_{\theta_{\text{old}}}(q))\mathbb{E}_{o\sim\pi_{\theta(.|q)}}\mathbbm{1}_{r(q,o)=1}-\omega^{-}_{\varepsilon}(p_{\theta_{\text{old}}}(q))\mathbb{E}_{o\sim\pi_{\theta}(.|q)}\mathbbm{1}_{r(q,o)=0}\Big{\}}-\beta\mathrm{KL}(\pi_{\theta}||\pi_{\mathrm{ref}}),$ (5)
Note that the formulation (GRPO-No-Clip) is the one currently implemented in HuggingFace TRL library (von Werra et al., 2020), hence we will focus in what follows on this version.
Algorithm 1 Iterative GRPO with verifiable rewards, modified from (Shao et al., 2024)
1:Input initial policy model $\pi_{\theta_{\text{init}}}$; verifiable reward $r$; task prompts $\mathcal{D}$; hyperparameters $\epsilon$, $\beta$, $\mu$
2:policy model $\pi_{\theta}\leftarrow\pi_{\theta_{\text{init}}}$
3:for $n=1,\dots,M$ do
4:    Sample a batch $\mathcal{D}_{b}$ from $\rho_{\mathcal{Q}}$
5:    Update the old policy model $\pi_{\theta_{\textrm{old}}}\leftarrow\pi_{\theta}$
6:    Sample $G$ outputs $\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\textrm{old}}}(\cdot\mid q)$ for each question $q\in\mathcal{D}_{b}$
7:    Compute rewards $\{r_{i}\}_{i=1}^{G}$ for each sampled output $o_{i}$ by running verifiable reward $r$
8:    Compute $A(q,o_{i})$ using equation (3), where $p=p_{\theta_{\text{old}}}(q)=\frac{1}{G}\sum_{i=1}^{G}\mathbbm{1}_{r(q,o_{i})=1}$
9:    for GRPO iteration = 1, …, $\mu$ do
10:         Update the policy model $\pi_{\theta}$ by maximizing the GRPO objective with gradient ascent (Equation (GRPO-No-Clip))
11:Output $\pi_{\theta}$
2.4. GRPO Iterations
Algorithm 1 summarizes GRPO iterations (Stabilized and no clipping). We see that GRPO iterations can be written as a sequence optimization we denote by $\pi_{\theta_{n}}$, the policy at iteration $n$. We see that GRPO iterations can be written for $n\geq 1$:
$\theta_{n}=\operatornamewithlimits{arg\,max}_{\theta}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\Big{\{}\omega^{+}_{\varepsilon}(p_{\theta_{n-1}}(q))\mathbb{E}_{o\sim\pi_{\theta(.|q)}}\mathbbm{1}_{r(q,o)=1}-\omega^{-}_{\varepsilon}(p_{\theta_{n-1}}(q))\mathbb{E}_{o\sim\pi_{\theta}(.|q)}\mathbbm{1}_{r(q,o)=0}\Big{\}}-\beta\mathrm{KL}(\pi_{\theta}||\pi_{\mathrm{ref}}),$ (6)
Note that in Algorithm 1, expectations are estimated using importance sampling from $\pi_{\theta_{n-1}}$, and each maximization problem is solved via gradient for $\mu$ steps.
In the following we will replace the maximization on the parameter space of the policy by maximizing over the space of policies (i.e optimization on the probability space) in order to analyze the dynamics of GRPO iterations as follows, for $n\geq 1$:
$\pi_{n}=\operatornamewithlimits{arg\,max}_{\pi}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\Big{\{}\omega^{+}_{\varepsilon}\left(p_{n-1}(q)\right)\mathbb{E}_{o\sim\pi(.|q)}\mathbbm{1}_{r(q,o)=1}-\omega^{-}_{\varepsilon}(p_{n-1}(q))\mathbb{E}_{o\sim\pi(.|q)}\mathbbm{1}_{r(q,o)=0}\Big{\}}-\beta\mathrm{KL}(\pi||\pi_{\mathrm{ref}}),$ (GRPO Iterations)
where $p_{n-1}(q)$ is the probability of success of the policy $\pi_{n-1}(\cdot|q)$:
$p_{n-1}(q)=\mathbb{E}_{o\sim\pi_{n-1}(.|q)}\mathbbm{1}_{r(q,o)=1}$ (7)
and the weights $\omega^{+}_{\varepsilon}$ and $\omega^{-}_{\varepsilon}$ are given in Equation (4). We assume all throughout the paper that
$\pi_{0}=\pi_{\mathrm{ref}}.$
Note that moving the optimization from a parametric space to the probability space can be seen as assuming that the hypothesis class of the parametric policies is large enough to represent all policies.
Note that in GRPO iterations the policy at iteration $n$ depends upon the policy $\pi_{n-1}$ via the probability of success $p_{n-1}$, as well on the reference policy via the $\mathsf{KL}$ regularizer.
3. GRPO Dynamics: Fixed Point iteration for Probability of Success
Our goal in this Section is to analyze the dynamics of the GRPO iterations given in Equation (GRPO Iterations).
Theorem 1 (GRPO Policy Dynamic).
Optimal GRPO iterations policies solving Equation (GRPO Iterations) satisfy the following recursion, for $n\geq 1$:
$\pi_{n}(o|q)=\frac{1}{Z_{n-1}(q)}\pi_{\mathrm{ref}}(o|q)\exp\left(\frac{1}{\beta}\left(\omega^{+}_{\varepsilon}(p_{n-1}(q))\mathbbm{1}_{r(q,o)=1}-\omega^{-}_{\varepsilon}(p_{n-1}(q))\mathbbm{1}_{r(q,o)=0}\right)\right),$
where
$Z_{n-1}(q)=p_{\mathrm{ref}}(q)\exp\left(\frac{1}{\beta}\omega^{+}_{\varepsilon}(p_{n-1}(q))\right)+(1-p_{\mathrm{ref}}(q))\exp\left(-\frac{1}{\beta}\omega^{-}_{\varepsilon}(p_{n-1}(q))\right),$
where the weights $\omega^{+}_{\varepsilon}$ and $\omega^{-}_{\varepsilon}$ are given in Equation (4), the probability of success $p_{n-1}(q)$ of policy $\pi_{n-1}(\cdot|q)$ is given in Equation (7), and $p_{\mathrm{ref}}(q)$ is the probability of success of the reference policy $\pi_{\mathrm{ref}}(\cdot|q)$: $p_{\mathrm{ref}}(q)=\mathbb{E}_{o\sim\pi_{\mathrm{ref}}(\cdot|q)}\mathbbm{1}_{r(q,o)=1}$.
We turn now to the recursion satisfied by the probability of success $p_{n}(q)$ of the policy $\pi_{n}(\cdot|q)$, we have the following theorem that shows that this success probability satisfies a fixed point iteration:
Theorem 2 (GRPO’s Probability of Success Fixed Point Iteration).
Assume $p_{\mathrm{ref}}>0$, define for $\beta>0$:
$h_{\varepsilon,p_{\mathrm{ref}}}(p)=\frac{1}{1+\frac{1-p_{\mathrm{ref}}}{p_{\mathrm{ref}}}\exp\left(-\frac{1}{\beta}\frac{1}{\sqrt{p(1-p)+\varepsilon}}\right)}$
The probability of success along GRPO’s iteration satisfies the following fixed point iteration i.e we have almost surely for all $q$ for $n\geq 1$
$p_{n}(q)=h_{\varepsilon,p_{\mathrm{ref}}(q)}(p_{n-1}(q)),$ (8)
and $p_{0}(q)=p_{\mathrm{ref}}(q)$.
Remark 1 (Importance of $\varepsilon>0$).
Note if $\varepsilon=0$, $h_{\varepsilon,p_{\mathrm{ref}}}$ is no longer continuous at 00 and $1$ and we can no longer guarantee existence of fixed points on $[0,1]$.
Figure 2. Fixed points as function of $\beta$ and $p_{\mathrm{ref}}$ for $\varepsilon=1e^{-5}$.
We study in the following proposition propreties of the function $h_{\varepsilon,p_{\mathrm{ref}}}$:
Proposition 1 (Propreties of $h_{\varepsilon,p_{\mathrm{ref}}}$).
$h_{\varepsilon,p_{\mathrm{ref}}}$ satisfies the following propreties:
  • Existence of fixed points: $h_{\varepsilon,p_{\mathrm{ref}}}$ is continuous $[0,1]$ to $[0,1]$ and hence admits at least a fixed point $p^{*}$ in $[0,1]$ (no guarantees for a unique fixed point)
  • Monotonicity:
    $h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}(p)=-h_{\varepsilon,p_{\mathrm{ref}}}(p)(1-h_{\varepsilon,p_{\mathrm{ref}}}(p))\frac{1-2p}{2\beta\,[p(1-p)+\varepsilon]^{3/2}}$
    • if $p<\frac{1}{2}$, $h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}(p)<0$ and $h_{\varepsilon,p_{\mathrm{ref}}}(p)$ is decreasing
    • if $p>\frac{1}{2}$ $h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}(p)>0$ and $h_{\varepsilon,p_{\mathrm{ref}}}(p)$ is increasing
    • if $p=\frac{1}{2}$ $h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}(p)=0$ and $p=\frac{1}{2}$ achieves its minimum
We drop in the sequel $q$, when referring to the sequence $p_{n}(q)$, and write for short $p_{n}$ (the reader is referred to Remark 2 for a discussion). If the sequence defined in GRPO’s probability of success iteration (8) converges we have therefore by continuity of $h_{\varepsilon,p_{\mathrm{ref}}}$:
$p_{\infty}=\lim_{n\to\infty}p_{n}=\lim_{n\to\infty}h_{\varepsilon,p_{\mathrm{ref}}}(p_{n-1})=h_{\varepsilon,p_{\mathrm{ref}}}(\lim_{n\to\infty}p_{n-1})=h_{\varepsilon,p_{\mathrm{ref}}}(p_{\infty}),$
and hence $p_{\infty}=h_{\varepsilon,p_{\mathrm{ref}}}(p_{\infty})$, and the limit point probability of success of GRPO $p_{\infty}=p^{*}$ is a fixed point of $h_{\varepsilon,p}$ (fixed points exist by virtue of proposition 1). Note that the fixed point $p^{*}$ is indeed function of $q$, and this dependency in $h_{\varepsilon,p_{\mathrm{ref}}}$ is via $p_{\mathrm{ref}}(q)$.
We see in Figure 2 various plots of the function $h_{\varepsilon,p_{\mathrm{ref}}}$ for different values of $\beta$ and initialization $p_{\mathrm{ref}}$, as well as the plot of the function $y=p$. Fixed points correspond to the intersections of this line with the curve of $h_{\varepsilon,p_{\mathrm{ref}}}$. We see that the fixed points are not unique in general, and $p^{*}=1$ is almost always a fixed point.
4. GRPO: Fixed Point Iteration Convergence and Probability of Success Amplification
In this Section we answer the following two questions:
  1. (1)
    Under which conditions on $\beta$ and $p_{\mathrm{ref}}$ do fixed points of GRPO iterations, $p^{*}$, lead to a probability of success $p^{*}$ that is higher than the reference initialization $p_{0}=p_{\mathrm{ref}}$?
  2. (2)
    Under which conditions do we have local convergence of the GRPO’s probability of success sequence given in (8) to a fixed point $p^{*}$ of $h_{\varepsilon,p_{\mathrm{ref}}}$?
Theorem 3 (GRPO amplifies the probability of success).
Let $10<p_{\mathrm{ref}}<$ 1. Let $p^{*}$ be a fixed point of $h_{\varepsilon,p_{\mathrm{ref}}}$ we have $p^{*}>p_{\mathrm{ref}}$, if:
  1. (1)
    $p_{\mathrm{ref}}\leq\frac{1}{2}$ for all $\beta>0$.
  2. (2)
    $p_{\mathrm{ref}}>\frac{1}{2}$ and $\beta\cosh^{2}\left(\frac{1}{2\beta}\frac{1}{\sqrt{\frac{1}{4}+\varepsilon}}\right)\geq\frac{p_{\mathrm{ref}}(1-p_{\mathrm{ref}})(2p_{\mathrm{ref}}-1)}{2[p_{\mathrm{ref}}(1-p_{\mathrm{ref}})+\varepsilon]^{3/2}}.$
Figure 3. Condition for probability amplification on $\beta$ is mostly satisified only for high $p$ and small $\beta$ (Blue area)
We see from Theorem 3 that the fixed point $p^{*}$ of the GRPO iteration leads to an amplification of the probability of success of the reference model $p_{\mathrm{ref}}$. We see from Figure 3 that the condition on $\beta$ is rather mild, and it implies that for high $p_{\mathrm{ref}}$, $\beta$ needs to be selected not too small.
We now turn to the second question regarding the convergence of the GRPO sequence given in (8) to a fixed point $p^{*}$ of $h_{\varepsilon,p_{\mathrm{ref}}}$. Given the properties of $h_{\varepsilon,p_{\mathrm{ref}}}$, we can characterize the limit point of the GRPO iteration as $n\to\infty$ as follows, as a consequence of the local Banach fixed-point theorem:
Theorem 4 (Local Fixed Point Convergence).
Let $p^{*}$ be a fixed point of $h_{\varepsilon,p_{\mathrm{ref}}}$and assume that have $|h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}(p^{*})|<1$.Given that $h_{\varepsilon,p_{\mathrm{ref}}}$ and $h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}$ are continuous in $[0,1]$, then there exists $\delta>0$ such the iteration $p_{n}=h_{\varepsilon,p_{\mathrm{ref}}}(p_{n-1})$ converges to $p^{*}$, if $p_{0}=p_{\mathrm{ref}}\in[p^{*}-\delta,p^{*}+\delta]$. In other words under this condition we have:
$\lim_{n\to\infty}p_{n}=p^{*}.$
Lemma 1.
Let $p^{*}$ be a fixed point: $p^{*}=h_{\varepsilon,p_{\mathrm{ref}}}(p^{*})$, then we have:
$h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}(p^{*})$ $=-h_{\varepsilon,p_{\mathrm{ref}}}(p^{*})(1-h_{\varepsilon,p_{\mathrm{ref}}}(p^{*}))\frac{1-2p^{*}}{2\beta\,[p^{*}(1-p^{*})+\varepsilon]^{3/2}}$
$=p^{*}(1-p^{*})\frac{2p^{*}-1}{2\beta\,[p^{*}(1-p^{*})+\varepsilon]^{3/2}}$
One condition for local convergence is therefore to have:
$|h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}(p^{*})|=p^{*}(1-p^{*})\frac{|2p^{*}-1|}{2\beta\,[p^{*}(1-p^{*})+\varepsilon]^{3/2}}<1$
which is satisfied if :
$\beta>\mathcal{B}(p^{*})=p^{*}(1-p^{*})\frac{|2p^{*}-1|}{2[p^{*}(1-p^{*})+\varepsilon]^{3/2}}.$
Figure 4. Lower bound on $\beta$ to ensure local convergence of GRPO fixed point iteration.
We see from Figure 4 the lower bound on $\beta$ required to ensure local convergence of GRPO iterations to a fixed point $p^{*}$. Figure 5 shows iteration (8) as a function of $n$ for different values of $\beta$ and $p_{\mathrm{ref}}$. We see that in most cases, there is a sharp transition where we observe fast convergence to $1$ or to a fixed point $p^{*}$. For $\beta=5$ and $p_{\mathrm{ref}}=0.001$, we see a divergent behavior.
Figure 5. GRPO Recursion and convergence to fixed points of $h_{\varepsilon}$, for $\varepsilon=1e^{-5}$
Remark 2.
Note that both conditions on $\beta$ are stated conditionally on a prompt $q$, to obtain results uniformly on $q$ we need to take $\sup$ on $q$ in all lower bounds.
5. Back to Parametric GRPO Iterations
Let $\tilde{\pi}_{n}=\pi_{\theta_{n}}$, the sequence of parametric policies solutions of problem (6) produced by gradient descent for example as in Algorithm 1. We make the following assumption on the total variation distance $\mathrm{TV}$ between these parametric policies and the non-parametric GRPO policies $\pi_{n}$ given in Theorem 1. We show in this Section if we have approximate policies we can have still asymptotic convergence.
Assumption 1.
We assume $\tilde{\pi}_{0}=\pi_{0}=\pi_{\mathrm{ref}}$ and assume for all $n\geq 1$, there exists $\delta_{n}\geq 0$ such that:
$\mathrm{TV}(\tilde{\pi}_{n}||\pi_{n})\leq\mathrm{TV}(\tilde{\pi}_{n-1}||\pi_{n-1})+\delta_{n},$
such that there exists $\delta^{*}\in[0,1)$ such that $\sum_{i=1}^{n}\delta_{i}\to\delta^{*}$ as $n\to\infty$.
We have the following theorem:
Theorem 5.
Under Assumption 1 and assuming that $p_{n}$ converges to $p^{*}$ the fixed point of $h_{\varepsilon,p_{\mathrm{ref}}}$. Let $\tilde{p}_{n}$ the probability of success of the policy $\tilde{\pi}$ we have:
$\lim_{n\to\infty}|\tilde{p}_{n}-p^{*}|\leq 2\delta^{*}.$
In the case $\delta^{*}=0$, we have convergence to the fixed point.
In Assumption 1 $\delta_{n}$ represent statistical, approximation and optimization errors. We see from Theorem 5, that as long these error remain small, the probability of success of GRPO parametric policy (estimated from samples and optimized for instance with gradient descent) remains close to the fixed point probability success $p^{*}$.
6. Conclusion
In conclusion, we have shown that GRPO with verifiable rewards can be viewed as an adaptive weighted contrastive loss (Section 2). We derived a closed-form recursion for the optimal policy, expressed in terms of the reference and old policies, and the probability of success (Section 3). This leads to a fixed-point equation, with GRPO’s probability of success converging to a fixed-point solution under mild assumptions (Section 4). Moreover, we proved that GRPO amplifies the probability of success compared to the reference model (Section 4). Finally, we showed that for approximate policies, the probability of success remains close to the fixed-point value as long as approximation statistical and optimization errors are small (Section 5).
Appendix A Proofs of Section 3
Proof of Theorem 1.
The objective in Equation (GRPO Iterations) is concave and hence setting the first order optimality conditions (See for example (Mroueh, 2024) ) we obtain:
$\pi_{n}(o|q)=\frac{1}{Z_{n-1}(q)}\pi_{\mathrm{ref}}(o|q)\exp\left(\frac{1}{\beta}\left(\omega^{+}_{\varepsilon}(p_{n-1}(q))\mathbbm{1}_{r(q,o)=1}-\omega^{-}_{\varepsilon}(p_{n-1}(q))\mathbbm{1}_{r(q,o)=0}\right)\right),$
where
[…]
$=\mathbb{E}_{o\sim\pi_{\mathrm{ref}}(\cdot|q)}\mathbbm{1}_{r(q,o)=1}\exp\left(\frac{1}{\beta}\left(\omega^{+}_{\varepsilon}(p_{n-1}(q))\mathbbm{1}_{r(q,o)=1}-\omega^{-}_{\varepsilon}(p_{n-1}(q))\mathbbm{1}_{r(q,o)=0}\right)\right)$
$+\mathbb{E}_{o\sim\pi_{\mathrm{ref}}(\cdot|q)}\mathbbm{1}_{r(q,o)=0}\exp\left(\frac{1}{\beta}\left(\omega^{+}_{\varepsilon}(p_{n-1}(q))\mathbbm{1}_{r(q,o)=1}-\omega^{-}_{\varepsilon}(p_{n-1}(q))\mathbbm{1}_{r(q,o)=0}\right)\right)$
$=\exp\left(\frac{1}{\beta}\omega^{+}_{\varepsilon}(p_{n-1}(q))\right)\mathbb{E}_{o\sim\pi_{\mathrm{ref}}(\cdot|q)}\mathbbm{1}_{r(q,o)=1}+\exp\left(-\frac{1}{\beta}\omega^{-}_{\varepsilon}(p_{n-1}(q))\right)\mathbb{E}_{o\sim\pi_{\mathrm{ref}}(\cdot|q)}\mathbbm{1}_{r(q,o)=0}$
$=p_{\mathrm{ref}}(q)\exp\left(\frac{1}{\beta}\omega^{+}_{\varepsilon}(p_{n-1}(q))\right)+(1-p_{\mathrm{ref}}(q))\exp\left(-\frac{1}{\beta}\omega^{-}_{\varepsilon}(p_{n-1}(q))\right),$
where
$p_{\mathrm{ref}}(q)=p_{0}(q)=\mathbb{E}_{o\sim\pi_{\mathrm{ref}}(\cdot|q)}\mathbbm{1}_{r(q,o)=1}.$
Proof of Theorem 2.
Replacing $\pi_{n}(\cdot|q)$ by its expression from Theorem 1 we have:
$p_{n}(q)$ $=\mathbb{E}_{o\sim\pi_{n}(.|q)}\mathbbm{1}_{r(q,o)=1}$
[…]
$=\frac{1}{Z_{n-1}(q)}\exp\left(\frac{1}{\beta}\omega^{+}_{\varepsilon}(p_{n-1}(q))\right)\mathbb{E}_{\pi_{\mathrm{ref}}}\mathbbm{1}_{r(q,o)=1}$
$=\frac{p_{\mathrm{ref}}(q)\exp\left(\frac{1}{\beta}\omega^{+}_{\varepsilon}(p_{n-1}(q))\right)}{Z_{n-1}(q)}$
$=\frac{p_{\mathrm{ref}}(q)\exp\left(\frac{1}{\beta}\omega^{+}_{\varepsilon}(p_{n-1}(q))\right)}{p_{\mathrm{ref}}(q)\exp\left(\frac{1}{\beta}\omega^{+}_{\varepsilon}(p_{n-1}(q))\right)+(1-p_{\mathrm{ref}}(q))\exp\left(-\frac{1}{\beta}\omega^{-}_{\varepsilon}(p_{n-1}(q))\right)}$
Replacing the weights expressions from Equations (4) we obtain:
$p_{n}(q)=\frac{p_{\mathrm{ref}}\exp\left(\frac{1}{\beta}\left(\frac{1-p_{n-1}(q)}{\sqrt{p_{n-1}(q)(1-p_{n-1}(q))+\varepsilon}}\right)\right)}{p_{\mathrm{ref}}\exp\frac{1}{\beta}\left(\frac{1-p_{n-1}(1)}{\sqrt{p_{n-1}(q)(1-p_{n-1}(q))+\varepsilon}}\right)+(1-p_{\mathrm{ref}})\exp\frac{1}{\beta}\left(-\frac{p_{n-1}(q)}{\sqrt{p_{n-1}(q)(1-p_{n-1}(q))+\varepsilon}}\right)}$ (9)
Define
$h_{\varepsilon,p_{\mathrm{ref}}}(p)=\frac{p_{\mathrm{ref}}\exp\left(\frac{1}{\beta}\left(\frac{1-p}{\sqrt{p(1-p)+\varepsilon}}\right)\right)}{p_{\mathrm{ref}}\exp\frac{1}{\beta}\left(\frac{1-p}{\sqrt{p(1-p)+\varepsilon}}\right)+(1-p_{\mathrm{ref}})\exp\frac{1}{\beta}\left(-\frac{p}{\sqrt{p(1-p)+\varepsilon}}\right)}$
We see therefore that GRPO’s probability of success satisfies the following iteration :
$p_{n}(q)=h_{\varepsilon,p_{\mathrm{ref}}}(p_{n-1}(q)).$
We assume here that $10<p_{\mathrm{ref}}<$ 1.We can simplify $h_{\varepsilon}(p)$ as follows:
$h_{\varepsilon,p_{\mathrm{ref}}}(p)$ $=\frac{1}{1+\frac{1-p_{\mathrm{ref}}}{p_{\mathrm{ref}}}\exp\frac{1}{\beta}\left(\frac{-p}{\sqrt{p(1-p)+\varepsilon}}-\frac{1-p}{\sqrt{p(1-p)+\varepsilon}}\right)}$
$=\frac{1}{1+\frac{1-p_{\mathrm{ref}}}{p_{\mathrm{ref}}}\exp\left(-\frac{1}{\beta}\frac{1}{\sqrt{p(1-p)+\varepsilon}}\right)}.$
Proof of Proposition 1.
Existence of fixed points For $\varepsilon>0$ $h_{\varepsilon,p_{\mathrm{ref}}}$ is continuous function from $[0,1]$ to $[0,1]$ and hence by Brouwer’s Fixed Point Theorem at least a fixed point $p^{*}$ exists in $[0,1]$, i.e $\exists p^{*}\in[0,1]$ such that $p^{*}=h_{\varepsilon,p_{\mathrm{ref}}}(p^{*})$.
Monotonicity Let $\sigma(z)=\frac{1}{1+\exp(-z)}$ and let $A=\frac{1-p_{\mathrm{ref}}}{p_{\mathrm{ref}}}$ and $B(p)=\frac{1}{\beta}\frac{1}{\sqrt{p(1-p)+\varepsilon}}$ hence we have:
$h_{\varepsilon,p_{\mathrm{ref}}}(p)=\sigma\left(z(p)\right)$
where
$z(p)=-\log(A)+B(p)$
we have
$z^{\prime}(p)=B^{\prime}(p)=-\frac{1-2p}{2\beta\,[p(1-p)+\varepsilon]^{3/2}}$
Let us compute the derivative :
$h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}(p)$ $=\sigma(z(p))(1-\sigma(z(p)))z^{\prime}(p)$
$=-\sigma(z(p))(1-\sigma(z(p)))\frac{1-2p}{2\beta\,[p(1-p)+\varepsilon]^{3/2}}$
  • if $p<\frac{1}{2}$, $h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}(p)<0$ and $h_{\varepsilon,p_{\mathrm{ref}}}$ is decreasing
  • if $p>\frac{1}{2}$ $h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}(p)>0$ and $h_{\varepsilon,p_{\mathrm{ref}}}$ is increasing
  • if $p=\frac{1}{2}$ $h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}(p)=0$
Appendix B Proofs of Section 4
Proof of Theorem 3.
We claim that any fixed point $p^{*}$ of $h_{\varepsilon}$ satisfies
$p^{*}>p_{\mathrm{ref}}.$
This is seen by considering the function
$f(p)=h_{\varepsilon}(p)-p,$
whose zeros correspond to fixed points.
Lemma 2.
For all $\beta,\varepsilon>0$, we have $f(p_{\mathrm{ref}})>0$.
Proof.
We have for all $\beta,\varepsilon>0$ $\exp\!\left(-\frac{1}{\beta}\frac{1}{\sqrt{p_{\mathrm{ref}}(1-p_{\mathrm{ref}})+\varepsilon}}\right)\leq 1$.
$f(p_{\mathrm{ref}})-p_{\mathrm{ref}}$ $=\frac{1}{1+\frac{1-p_{\mathrm{ref}}}{p_{\mathrm{ref}}}\,\exp\!\left(-\frac{1}{\beta}\frac{1}{\sqrt{p_{\mathrm{ref}}(1-p_{\mathrm{ref}})+\varepsilon}}\right)}-p_{\mathrm{ref}}$
$>\frac{1}{1+\frac{1-p_{\mathrm{ref}}}{p_{\mathrm{ref}}}}-p_{\mathrm{ref}}$
$=p_{\mathrm{ref}}-p_{\mathrm{ref}}$
$=0.$
Let us compute the derivative of $f$ we have:
$f^{\prime}(p)=h_{\varepsilon}^{\prime}(p)-1=-\sigma(z(p))(1-\sigma(z(p)))\frac{1-2p}{2\beta\,[p(1-p)+\varepsilon]^{3/2}}-1$
If $0<p<\frac{1}{2}$ we have $f^{\prime}(p)<0$ and $f$ is decreasing. If $p_{\mathrm{ref}}<\frac{1}{2}$ We have $f(0)=1>0$ and $f(p_{\mathrm{ref}})>0$, since $f$ decreasing in $[0,p_{\mathrm{ref}}]\subset(0,\frac{1}{2})$ and is strictly positive in this interval. $f(p)=0$ will be in $(p_{\mathrm{ref}},1]$ and hence $p^{*}>p_{\mathrm{ref}}$.
Now we know that the minimum of $h_{\varepsilon}(p)$ is achieved for $p=\frac{1}{2}$. We have
$h_{\varepsilon}\left(\frac{1}{2}\right)=\frac{1}{1+\frac{1-p_{\mathrm{ref}}}{p_{\mathrm{ref}}}\,\exp\!\left(-\frac{1}{\beta}\frac{1}{\sqrt{\frac{1}{4}+\varepsilon}}\right)}>p_{\mathrm{ref}}$
For $p_{\mathrm{ref}}>\frac{1}{2}$, we have $h_{\varepsilon}(p_{\mathrm{ref}})>h_{\varepsilon}\left(\frac{1}{2}\right)>p_{\mathrm{ref}}>\frac{1}{2}$, since $h_{\varepsilon}(p)$ is increasing on $[\frac{1}{2},1]$. It follows that $f(\frac{1}{2})>0$ and $f(p_{\mathrm{ref}})>0$
We would like to to find a condition on $\beta$ for $p_{\mathrm{ref}}>\frac{1}{2}$ such that for $p\in(\frac{1}{2},p_{\mathrm{ref}}]$:
$f^{\prime}(p)=\sigma(z(p))(1-\sigma(z(p)))\frac{2p-1}{2\beta\,[p(1-p)+\varepsilon]^{3/2}}-1<0$
where $z(p)=\underbrace{-\log\left(\frac{1-p_{\mathrm{ref}}}{p_{\mathrm{ref}}}\right)}_{a(p_{\mathrm{ref}})}+\frac{1}{\beta}\underbrace{\frac{1}{\sqrt{p(1-p)+\varepsilon}}}_{b(p)}$. Note that for $p_{\mathrm{ref}}>\frac{1}{2},a(p_{\mathrm{ref}})>0$ and $\beta(p)>0$. Using Lemma 2 we have for $p\in(\frac{1}{2},p_{\mathrm{ref}}]$:
$f^{\prime}(p)$ $=\sigma(z(p))(1-\sigma(z(p)))\frac{2p-1}{2\beta\,[p(1-p)+\varepsilon]^{3/2}}-1$
$=\frac{1}{4\cosh^{2}(\frac{z(p)}{2})}\frac{2p-1}{2\beta\,[p(1-p)+\varepsilon]^{3/2}}-1$ (Using Lemma 2 (1))
$=\frac{1}{4\cosh^{2}\left(\frac{a(p_{\mathrm{ref}})+\frac{1}{\beta}b(p)}{2}\right)}\frac{2p-1}{2\beta\,[p(1-p)+\varepsilon]^{3/2}}-1$
Note that $p\to\frac{2p-1}{2\beta\,[p(1-p)+\varepsilon]^{3/2}}$ is increasing in $(\frac{1}{2},p_{\mathrm{ref}}]$ and hence we have for $p\in(\frac{1}{2},p_{\mathrm{ref}}]$
$0<\frac{2p-1}{2\beta\,[p(1-p)+\varepsilon]^{3/2}}\leq\frac{2p_{\mathrm{ref}}-1}{2\beta\,[p_{\mathrm{ref}}(1-p_{\mathrm{ref}})+\varepsilon]^{3/2}}$
On the other hand by Lemma 2 (2) we have:
$\frac{1}{4\cosh^{2}\left(\frac{a(p_{\mathrm{ref}})+\frac{1}{\beta}b(p)}{2}\right)}$ $\leq\frac{1}{4}\frac{1}{\cosh^{2}(\frac{a(p_{\mathrm{ref}})}{2})\cosh^{2}(\frac{b(p)}{2\beta})}$
$=\frac{1}{4}\frac{1}{\cosh^{2}(\frac{a(p_{\mathrm{ref}})}{2})\cosh^{2}(\frac{b(p)}{2\beta})}$
$=\frac{4p_{\mathrm{ref}}(1-p_{\mathrm{ref}})}{4}\frac{1}{\cosh^{2}(\frac{b(p)}{2\beta})}$ Using Lemma 3 (2)
Now let us study the monotonicity of $\xi(p)=\cosh^{2}(\frac{1}{2\beta}b(p))$, we have $\xi^{\prime}(p)=\frac{1}{2\beta}b^{\prime}(p)(\cosh^{2}(\frac{1}{2\beta}b(p)))^{\prime}=\frac{1}{2\beta}b^{\prime}(p)\sinh(2\frac{1}{2\beta}b(p)))=\frac{1}{4\beta}\frac{(2p-1)}{(p(1-p)+\varepsilon)^{\frac{3}{2}}}\sinh(\frac{1}{\beta}b(p)))$ Hence we see that for $p\in(\frac{1}{2},p_{\mathrm{ref}}]$ $\xi^{\prime}(p)>0$ and $\xi$ is increasing on $(\frac{1}{2},p_{\mathrm{ref}}]$, and hence we have:
$\xi\left(\frac{1}{2}\right)\leq\xi(p)\leq\xi(p_{\mathrm{ref}})$
hence we have:
$\frac{1}{\xi(p_{\mathrm{ref}})}\leq\frac{1}{\xi(p)}\leq\frac{1}{\xi\left(\frac{1}{2}\right)}=\frac{1}{\cosh^{2}(\frac{1}{2\beta}\frac{1}{\sqrt{\frac{1}{4}+\varepsilon}})}$
It follows that we have:
$0<\frac{1}{4\cosh^{2}\left(\frac{a(p_{\mathrm{ref}})+\frac{1}{\beta}b(p)}{2}\right)}$ $\leq p_{\mathrm{ref}}(1-p_{\mathrm{ref}})\frac{1}{\cosh^{2}(\frac{1}{2\beta}\frac{1}{\sqrt{\frac{1}{4}+\varepsilon}})}$
Hence we have for $p\in(0.5,p_{\mathrm{ref}}]$
$f^{\prime}(p)$ $\leq\frac{2p_{\mathrm{ref}}-1}{2\beta\,[p_{\mathrm{ref}}(1-p_{\mathrm{ref}})+\varepsilon]^{3/2}}p_{\mathrm{ref}}(1-p_{\mathrm{ref}})\frac{1}{\cosh^{2}(\frac{1}{2\beta}\frac{1}{\sqrt{\frac{1}{4}+\varepsilon}})}-1$
$=\frac{p_{\mathrm{ref}}(1-p_{\mathrm{ref}})(2p_{\mathrm{ref}}-1)}{2[p_{\mathrm{ref}}(1-p_{\mathrm{ref}})+\varepsilon]^{3/2}}\frac{1}{\beta\cosh^{2}(\frac{1}{2\beta}\frac{1}{\sqrt{\frac{1}{4}+\varepsilon}})}-1$
To have $f^{\prime}(p)\leq 0$ we need to have:
$\beta\cosh^{2}\left(\frac{1}{2\beta}\frac{1}{\sqrt{\frac{1}{4}+\varepsilon}}\right)\geq\frac{p_{\mathrm{ref}}(1-p_{\mathrm{ref}})(2p_{\mathrm{ref}}-1)}{2[p_{\mathrm{ref}}(1-p_{\mathrm{ref}})+\varepsilon]^{3/2}}$
Lemma 3.
We have:
  1. (1)
    $\sigma(z)(1-\sigma(z))=\frac{1}{4\cosh^{2}(z/2)}$
  2. (2)
    For $z,a>0$, $\sigma(z+a)(1-\sigma(z+a))\leq\frac{1}{4\cosh^{2}(z/2)\cosh^{2}(a/2)}$
Proof.
(1) Note that $\cosh(x)=\frac{1}{2}(e^{x}+e^{-x})$ and $\cosh^{2}(x)=\frac{1}{4}(e^{2x}+e^{-2x}+2)=\frac{1}{4}e^{2x}(1+e^{-4x}+2e^{-2x})=\frac{1}{4}\left(1+e^{-2x}\right)^{2}e^{2x}.$ Recall that :
$\sigma(z)(1-\sigma(z))$ $=\frac{e^{-z}}{(1-e^{-z})^{2}}$
It follows that $\cosh^{2}(x)=\frac{1}{4}\sigma(2z)(1-\sigma(2z))$, which gives us finally:
$\sigma(z)(1-\sigma(z))=\frac{1}{4\cosh^{2}(\frac{z}{2})}$
(2) Note that we have for $a,z>0$
$\cosh\left(\frac{z+a}{2}\right)$ $=\cosh\left(\frac{z}{2}\right)\cosh\left(\frac{a}{2}\right)+\sinh\left(\frac{z}{2}\right)\sinh\left(\frac{a}{2}\right)\geq\cosh\left(\frac{z}{2}\right)\cosh\left(\frac{a}{2}\right),$
since $\sinh\left(\frac{z}{2}\right)\geq 0$ $\sinh\left(\frac{a}{2}\right)\geq 0$ for $a,z>0$. and hence we have:
$\sigma(z+a)(1-\sigma(z+a))=\frac{1}{4\cosh^{2}((z+a)/2)}\leq\frac{1}{4}\frac{1}{\cosh^{2}(\frac{a}{2})\cosh^{2}(\frac{z}{2})}.$
Lemma 4.
We have for $p_{\mathrm{ref}}>\frac{1}{2}$, $a(p_{\mathrm{ref}})=-\log(\frac{1-p_{\mathrm{ref}}}{p_{\mathrm{ref}}})>0$, and $\cosh^{2}(\frac{a(p_{\mathrm{ref}})}{2})=\frac{1}{4p_{\mathrm{ref}}(1-p_{\mathrm{ref}})}.$
Proof.
Let $x=\frac{a(p_{\mathrm{ref}})}{2}=\log(\sqrt{\frac{p_{\mathrm{ref}}}{1-p_{\mathrm{ref}}}})$ we have $\cosh(x)=\frac{1}{2}(e^{x}+e^{-x})=\frac{1}{4}(\frac{(1-p_{\mathrm{ref}})^{2}}{p_{\mathrm{ref}}^{2}}+\frac{p_{\mathrm{ref}}^{2}}{(1-p_{\mathrm{ref}})^{2}}+2)=\frac{1}{2}(\sqrt{\frac{p_{\mathrm{ref}}}{1-p_{\mathrm{ref}}}}+\sqrt{\frac{1-p_{\mathrm{ref}}}{p_{\mathrm{ref}}}})$. and hence:
$\cosh^{2}(x)$ $=\frac{1}{4}\left(\frac{p_{\mathrm{ref}}}{1-p_{\mathrm{ref}}}+\frac{1-p_{\mathrm{ref}}}{p_{\mathrm{ref}}}+2\right)$
$=\frac{1}{4}\frac{p_{\mathrm{ref}}^{2}+(1-p_{\mathrm{ref}})^{2}+2p_{\mathrm{ref}}(1-p_{\mathrm{ref}})}{p_{\mathrm{ref}}(1-p_{\mathrm{ref}})}$
$=\frac{1}{4}\frac{(p_{\mathrm{ref}}+1-p_{\mathrm{ref}})^{2}}{p_{\mathrm{ref}}(1-p_{\mathrm{ref}})}$
$=\frac{1}{4p_{\mathrm{ref}}(1-p_{\mathrm{ref}})}.$
Proof of Theorem 4.
This is a direct application of local Banach fixed point theorem:
Theorem 6 (Local Contraction Mapping for One-Dimensional Functions).
Let $f:\mathbb{R}\to\mathbb{R}$ be continuously differentiable, and suppose that $x^{*}\in\mathbb{R}$ is a fixed point of $f$ (i.e., $f(x^{*})=x^{*}$). Assume that $f^{\prime}$ is continuous and that
$\lvert f^{\prime}(x^{*})\rvert<1.$
Then, by the continuity of $f^{\prime}$, there exists a radius $r>0$ and a constant $k$ with
$\lvert f^{\prime}(x)\rvert\leq k<1\quad\text{for all }x\in[x^{*}-r,\,x^{*}+r].$
Consequently, $f$ is a contraction on the interval $I=[x^{*}-r,x^{*}+r]$, and for any initial guess $x_{0}\in I$, the iteration defined by
$x_{n+1}=f(x_{n})$
converges to the unique fixed point $x^{*}$ in $I$.
Appendix C Proofs of Section 5
Proof of Theorem 5.
Note that
$\mathrm{TV}(\tilde{\pi}||\pi)=\frac{1}{2}\sup_{||f||_{\infty}}\mathbb{E}_{\tilde{\pi}}f-\mathbb{E}_{{\pi}}f$
We have:
$|\tilde{p}_{n}-p_{n}|$ $=\Big{|}\mathbb{E}_{\tilde{\pi}_{n}}\mathbbm{1}_{r(q,o)=1}-\mathbb{E}_{\pi_{n}}\mathbbm{1}_{r(q,o)=1}\Big{|}$
$\leq 2\mathrm{TV}(\tilde{\pi}_{n}||\pi_{n})$
$\leq 2\sum_{i=1}^{n}\delta_{i}+\mathrm{TV}(\tilde{\pi}_{0},\pi_{0})$
$=2\sum_{i=1}^{n}\delta_{i}.$
Assume the sequence $p_{n}$ converges to $p^{*}$ the fixed point of $h_{\varepsilon,p_{\mathrm{ref}}}$. Under Assumption 1 we have :
$\lim_{n\to\infty}|\tilde{p}_{n}-p_{n}|\leq 2\lim_{n\to\infty}\sum_{i=1}^{n}\delta_{i}=2\delta^{*}$